sum of series 1/2+1/3+1/4+..........to n terms i.e.,1/2,1/3 ,1/4................are in HP

Grade:12

1 Answers

Shubham
6 years ago
1/a,1/b,1/c,.....are in HP then it`s reciprocal a,b,c,.....are in AP2,3,4,……are in APThen we put the formula for sum of n th termsSn=n/2[2a+(n-1)d] where ,a=2 and ,d=1Sn=n/2[4+(n-1)]